A married individual who is paid on a monthly basis, claims 5 federal withholding allowances, and earns $4,892 has $______ in federal income tax withholding when using the wage-bracket method.

(what is the amount of federal income tax withholding)
Answer #1

Monthly Wages = 4892

Since the wage amount is greater than 3593, hence wage bracket table will not be used.

We will use the percentage method to find the Federal Income Tax.

Monthly Wage = 4892

One Withholding Allowance for Monthly payroll = 350

Number of withholding allowance claim = 5

Total Withholding Allowance = 5*350 = 1750

Taxable Wages = 4892 - 1750 = 3142

From Table -

Federal Tax = 161.70 + (3142-2600)*12%

Federal Tax = 161.70 + 65.04

Federal Tax = $ 226.74